OpenStudy (anonymous):

What is the average rate of change from x = 1 to x = 3 for the function g(x) = ln (x + 2)? Round the answer to the nearest hundredth

OpenStudy (perl):

$$\Large{ \textrm{average rate of change} =\frac{g(3)-g(1)}{3-1} \\ =\frac{\ln(3+2)-\ln(1+2)}{3-1} }$$

OpenStudy (anonymous):

I'm not sure if my calculator did this correctly, but the answer I was given was 0.24

OpenStudy (anonymous):

Though, I typed it in a different way and was given the answer: 1.17

OpenStudy (anonymous):

Thank you so much! I typed it in incorrectly, I can see now.. I appreciate this so much! I hope you have a wonderful day!